# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v1

@@ 6.1. Experimental Setup
  • ReAcTable (Zhang et al., 2024). ReAcTable uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs. Once all relevant data is gathered, it asks the LLMs to predict the answer. We run the original code from (cod, 2024) and kept all settings as default.
Implementation Details. We mainly use DeepSeek (Guo et al., 2024) API [^2] https://platform.deepseek.com/usage model as the foundation model for each LLM-based agent. We also explore other proprietary LLMs, such as GPT-3.5 (Brown et al., 2020), using the GPT-3.5-turbo-0613 version, as adopted by (Wang et al., 2024b). To ensure reproducibility, we set the temperature parameter for all agents to 0.01.
@@ 6.2. Data Prep Method Comparison
Table 2. Comparison of AutoPrep with data prep methods for TQA (the best results are bolded, and the second-best results are underlined).
Method DeepSeek GPT3.5
WikiTQ TabFact WikiTQ TabFact
Binder $56.98$ $79.64$ $\underline{56.74}$ $\underline{79.17}$
ReAcTable $\underline{64.25}$- $51.80$ $72.80$
AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$ (87.01) $\mathbf{63.86}$ $\mathbf{84.29}$
Exp-1: Comparison of AutoPrep with other data prep methods for TQA. We conduct experiments to evaluate AutoPrep against current state-of-the-art data prep methods for TQA on two datasets: WikiTQ and TabFact. Specifically, we compare AutoPrep with Binder and ReAcTable, presenting the results in Table 2. Notably, due to the lack of public scripts for TabFact from ReAcTable (Zhang et al., 2024), we report results only for DeepSeek on WikiTQ and use GPT-3.5 results from their original paper.
The results clearly demonstrate that AutoPrep achieves state-of-the-art performance on both the WikiTQ and TabFact datasets when using DeepSeek and GPT-3.5 as the foundation model. In particular, AutoPrep outperforms Binder by 14.86% on WikiTQ and 8.50% on TabFact with DeepSeek. When using GPT-3.5, the performance improvements are 12.55% and 6.47%, respectively. Furthermore, when compared with ReAcTable, AutoPrep enhances performance by 1.87% on WikiTQ using DeepSeek. Remarkably, with GPT-3.5, AutoPrep surpasses ReAcTable by 23.28% and 15.78% on WikiTQ and TabFact, respectively.
These improvements are attributed to AutoPrep’s comprehensive data prep operations, which are not fully addressed by Binder and ReAcTable. While Binder integrates LLMs to generate new columns to supplement external knowledge, it neglects critical aspects such as column filering and value normalization. In contrast, ReAcTable employs the ReAct data processing paradigm but lacks a specific decomposition of data prep tasks. Our experimental results highlight the importance of thorough data prep in enhancing TQA performance, which is often a significant bottleneck for existing methods.
Additionally, as shown in Table 2, DeepSeek consistently outperforms GPT-3.5 while incurring lower API costs [^3] https://platform.deepseek.com/api-docs/zh-cn/quick_start/pricing/ https://openai.com/api/pricing/. Therefore, we have selected DeepSeek as our default foundation model for subsequent experiments.
6.3. Improvement of Data Prep for TQA
Table 3. Impact of data prep on TQA Methods.
Method  WikiTQ  TabFact
Zero-Shot QA $17.66$ $76.78$
    + AutoPrep $28.22_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 10.56}}$ $78.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 1.92}}$
Few-Shot QA $56.05$ $81.97$
    + AutoPrep $56.33_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.28}}$ $82.70_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 0.73}}$
Text-to-SQL $51.66$ $68.53$
    + AutoPrep $\mathbf{65.45_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 13.79}}}$ $\mathbf{86.41_{{\color[rgb]{0.1328125,0.6015625,0.328125}\uparrow 17.88}}}$
(a) Zero-Shot QA.
(b) Few-Shot QA.
(c) Text-to-SQL.
Figure 7. Performance on Small, Medium, Large tables from WikiTQ without/with data prep of AutoPrep.
Exp-2: Impact of data prep on TQA performance. We evaluate the performance of three baseline TQA methods, namely Zero-Shot QA, Few-Shot QA, and Text-to-SQL with and without the data prep provided by AutoPrep. The results are detailed in Table 3.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, Text-to-SQL experiences the most substantial enhancement, with performance increases of 26.69% on WikiTQ and 26.09% on TabFact. This significant improvement is attributed to Text-to-SQL’s sensitivity to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
Moreover, Zero-Shot and Few-Shot QA methods also show considerable performance gains. Zero-Shot QA sees an improvement of 59.80% on WikiTQ and 2.50% on TabFact, while Few-Shot QA improves by 0.50% on WikiTQ and 0.89% on TabFact. These improvements are largely due to AutoPrep’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
@@ 6.3. Improvement of Data Prep for TQA
To further validate the impact of data prep, we analyze performance across tables of varying sizes. We categorize the tables in WikiTQ and TabFact into small (fewer than 2048 tokens), medium (2048 to 4096 tokens), and large (more than 4096 tokens) categories. While all tables in TabFact are small, WikiTQ includes 4040 small tables, 200 medium tables, and 104 large tables. As shown in Figure 7, AutoPrep achieves significant gains of 5.88% and 23.67% on medium and large tables, respectively. In addition, Few-Shot QA demonstrates limited improvement for small tables, because it has already achieved a high accuracy.
These experimental results clearly show that the integration of AutoPrep enhances the performance of all TQA methods, which demonstrates the critical role of question-aware data preparation in optimizing TQA performance.
6.4. Evaluation on Planner Agent
Table 4. Comparison of different Methods for Planner and Programmer.
Agent Method  WikiTQ  TabFact
Planner Direct Prompting $58.17$ $83.35$
Chain-of-Clauses $65.45$ $86.41$
Programmer Code-generation $62.82$ $81.97$
Tool-augmented $65.45$ $86.41$
@@ 6.4. Evaluation on Planner Agent
Exp-3: How does our proposed Analysis Sketch method in suggesting high-level operations compared with Direct Prompting method? We change the high-level operation suggestion methods in Planner and evaluate the framework variants on WikiTQ and TabFact dataset. The results are reported in Table 4.
@@
As shown in the table, Analysis Sketch based method outperforms Direct Prompting method by 12.52% and 3.67%, which indicates the superiority of our proposed method in generating more accurate high-level operations. Moreover, we find that the performance improvement on WikiTQ is much more significant than that in TabFact. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using direct prompting. In Summary, the experimental results prove that our proposed Analysis Sketch method is more effective in suggesting high-level operations, especially for difficult TQA instances.
6.5. Evaluation on Programmer Agents
Exp-4: How does our proposed tool-augmented method in generating low-level operations compared with existing code generation methods? We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default. We evaluate the performance on WikiTQ and TabFact and report the results in Table 4.
As illustrated in Table 4, our proposed tool-augmented method achieves better performance compared with the existing code generation method by an improvement of 4.19% and 5.42%. Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. The above experimental results show the superiority of our method in both accuracy and cost.
Table 5. The average debugging round for Code-generation (Code) and Tool-augmented (Tool) Method.
Method  WikiTQ  TabFact
Code $1.35$ $1.28$
Tool $1.0$ $1.0$
Moreover, to compare the debugging method of these two methods, we record the average debugging rounds in Table 5. Our tool-augmented method conduct debugging process with fewer rounds, implying the effectiveness of our proposed debugging method for tool-augmented method.
6.6. Ablation Study
Table 6. Ablation Study of AutoPrep.
Method  WikiTQ  TabFact
AutoPrep $\mathbf{65.45}$ $\mathbf{86.41}$
- Filter $62.78$ $84.98$
  - Augmenter $61.79$ $85.67$
  - Normalizer $62.02$ $83.79$
Exp-5: What is the contribution of each agent in AutoPrep? We conduct ablation experiments to evaluate the contribution of each agent in AutoPrep. Specifically, we ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
The results are reported in Table 6. For WikiTQ datasets, without column augmentation, the performance drops the most (5.92%). For TabFact datasets, the Normalizer matters the most with a performance drop by 3.13%. This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while column type normalization is a primary issue in TabFact datasets. Moreover, for both datasets, each agent plays an essential role in data preparation for TQA, which brings performance improvement by at least 4.25% and 0.86%.
The experimental results have demonstrated the rationality of the multi-agent design in AutoPrep. By solving three types of data preparation tasks in TQA methods in a specialized manner, the performance will be improved.
6.7. Case Study
@@
Figure 8. Two case studies selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
@@
This section presents two illustrative examples selected from the WikiTQ dataset to qualitatively analyze effectiveness of AutoPrep.
As shown in Figure 8(a), to answer the question based on the original table, a Text-to-SQL method tries to filter all records with a score greater than 30. However, due to the limitation of SQL grammar, it can only use the operation Like to extract records where number 31, 32, …, or 39 is in values of column Score. This could miss out values with score greater than 40. Thus, it outputs error answer “3”. AutoPrep improves the performance of TQA methods by solving question-related data prep issues of the table in two stages. Specifically, based on the question in Figure 8(a), we first suggest two required high-level operations, i.e., a $\mathtt{Filter}$ to select the related column Score and a $\mathtt{Augment}$ to generate a new column from Score representing whether the team score is over 30. Then, we select appropriate functions from specific function pool and implement corresponding arguments to implement these high-level operations. Finally, we output a prepared table which only has columns related to the question. By passing the processed table and original question into the Text-to-SQL method, it generate a SQL that can easily extract the correct answer “4”.
Figure 8(b) illustrates a TQA instance with redundancy and inconsistency issues. To answer the question, it requires to first extract records with AirDate previous to “November 15, 2007” and then select records with RankOverall between 30 to 39. However, the values in column AirDate can not be directly assessed for chronological order through string comparison and RankOverall values are not in numerical type. Thus, although the generate SQL query is logically correct, it produces an wrong answer. AutoPrep suggests three high-level operations to solve this. Specifically, AutoPrep adopts an $\mathtt{Filter}$ to extract all question-related columns. And utilize two $\mathtt{Normalize}$ operations to format values of AirDate into sortable strings and cast RankOverall column to the integer type. We select three low-level operations to implement these high-level operations and get a new table with AirDate in %Y-%m-%d and RankOverall in integers. Then, we adopt the Text-to-SQL method on the prepared table, which generates a correct answer “3”.

### v2

@@ 6.1. Experimental Setup
ReAcTable (Zhang et al., 2024) uses the ReAct paradigm to extract relevant data from the table using Python or SQL code generated by LLMs. Once all relevant data is gathered, it asks the LLMs to predict the answer. We run the original code from (cod, 2024b) and keep all settings as default. Notice that the original code does not include prompts for TabFact, we generate it based the WikiTQ prompt.
Chain-of-Table (CoTable) (Wang et al., 2024b) enhances the table reasoning capabilities of LLMs by predefining several common atomic operations (including data prep operations) that can be dynamically selected by the LLM. These operations form an “operation chain” that represents the reasoning process over a table and can be executed either via Python code or by prompting the LLM. We implement CoTable using the original code from (cod, 2024a).
Table 2. Improvement of data prep for TQA (the best results are in bold and the second-best are underlined).
Method DeepSeek GPT3.5 Llama3 QWen2.5
WikiTQ TabFact WikiTQ TabFact WikiTQ TabFact WikiTQ TabFact
End2End 56.6556.65{56.65}56.65 81.7781.77{81.77}81.77 $52.56$ 71.5471.54{71.54}71.54 58.7258.72{58.72}58.72 81.2781.27{81.27}81.27 60.0160.01{60.01}60.01 81.1781.17{81.17}81.17
    + AutoPrep $63.14$ $\uparrow 6.49$ $82.11$ $\uparrow 0.34$ $61.21$ $\uparrow 8.65$ $71.79$ $\uparrow 0.25$ $61.23$ $\uparrow 2.51$ $84.19$ $\uparrow 2.92$ $63.42$ $\uparrow 3.41$ $83.05$ $\uparrow 1.88$
CoT $54.95$ 82.0282.02{82.02}82.02 53.4853.48{53.48}53.48 $65.37$ $40.75$ 80.9380.93{80.93}80.93 59.6759.67{59.67}59.67 82.3182.31{82.31}82.31
+ AutoPrep $61.12$ $\uparrow 6.17$ $82.26$ $\uparrow 0.24$ $60.01$ $\uparrow 6.53$ $74.36$ $\uparrow 8.99$ $56.01$ $\uparrow 15.26$ $83.65$ $\uparrow 2.72$ $62.02$ $\uparrow 2.35$ $\underline{85.67}$ $\uparrow 3.36$
NL2Py 59.3559.35{59.35}59.35 $68.13$ 53.5953.59{53.59}53.59 66.1566.15{66.15}66.15 $50.12$ $76.24$ $53.02$ $72.63$
+ AutoPrep $\underline{65.86}$ $\uparrow 6.51$ $\underline{87.35}$ $\uparrow 19.22$ $\underline{64.69}$ $\uparrow 11.1$ $\underline{84.83}$ $\uparrow 18.68$ $\underline{62.55}$ $\uparrow 12.43$ $\underline{85.42}$ $\uparrow 9.18$ $\underline{68.65}$ $\uparrow 15.63$ 85.72 $\uparrow 13.09$
NL2SQL $52.83$ $70.21$ $52.90$ $64.71$ 51.8051.80{51.80}51.80 $75.15$ $56.86$ $80.09$
+ AutoPrep 66.09 $\uparrow 13.26$ 87.85 $\uparrow 17.64$ 64.75 $\uparrow 11.85$ 84.19 $\uparrow 19.48$ 63.72 $\uparrow 11.92$ 85.72 $\uparrow 10.57$ 68.72 $\uparrow 11.86$ $85.33$ $\uparrow 5.24$
Evaluation Metrics. We adopt the evaluator in Binder (Cheng et al., 2023) to avoid the situation where program executions are likely to be semantically correct but fail to match the golden answer exactly. For example, when the SQL outputs 1/0 for yes/no questions, the Binder evaluator will conduct a pre-matching check and translate it to boolean values first. It will avoid mismatches of correct prediction and golden answer, such as pairs like 7.45 and 7.4499, 4 years and 4. Afterwards, we adopt accuracy as our evaluation metric.
Backbone LLMs. We evaluate our method using representative LLMs as backbones. For closed-source LLMs, we select DeepSeek (Guo et al., 2024) (DeepSeek-V2.5-Chat) and GPT3.5 (Brown et al., 2020) (GPT3.5-Turbo-0613). For open-source LLMs, we choose Llama3 (Dubey et al., 2024) (Llama-3.1-70B-Instruct) and QWen2.5 (Yang et al., 2024) (QWen2.5-72B-Instruct) for evaluation.
Experiment Settings. We provide detailed prompts of each component in AutoPrep in our technical report (tec, 2025) due to the space limit. For a fair comparison, we set the maximum token input of all methods as 8192. Moreover, we set the temperature parameter of all methods to 0.01 for reproducibility.
6.2. Improvement of Data Prep for TQA
Exp-1: Impact of question-aware data prep on TQA performance. We integrate AutoPrep into our four TQA baselines w/o data prep, in order to investigate the impact of question-aware data prep on TQA performance. The results are reported in Table 2.
As demonstrated, integrating AutoPrep significantly improves the performance of all evaluated methods. Notably, NL2SQL shows the most substantial gains, achieving an average accuracy improvement of 12.22 on WikiTQ and 13.23 on TabFact across all LLM backbones. Similarly, NL2Py also shows notable improvements in its performance, after being integrated with AutoPrep. This significant improvement is attributed to the sensitivity of NL2SQL and NL2Py to data incompleteness and inconsistency, which can lead to erroneous outcomes when performing operations on improperly formatted data. Thus, data prep operations, such as Augment and Normalize can solve these cases and improve the overall results.
Moreover, End2End and CoT methods also show considerable performance gains. For example, End2End shows an average accuracy gain of 5.27 and 1.35 on WikiTQ and TabFact respectively. These improvements are largely due to AutoPrep ’s filtering mechanism, which removes irrelevant columns, thereby simplifying the reasoning process for extracting answers directly from tables.
Finding 1: The integration of AutoPrep improves the accuracy of TQA methods w/o data prep by 12.22 points on WikiTQ and 13.23 points on TabFact, respectively.
Since “NL2SQL + AutoPrep” achieves the best accuracy in most cases, we take its results as default for further comparison.
@@ 6.3. Data Prep Method Comparison
Table 3. Experimental results of AutoPrep and TQA methods with data prep.
Method DeepSeek GPT3.5 Llama3 QWen2.5
 WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact  WikiTQ  TabFact
ICL-Prep 56.5456.54{56.54}56.54 $80.53$ $55.71$ $73.91$ $50.05$ $75.20$ $57.00$ $80.14$
Dater $48.32$ $83.05$ $52.81$ $72.08$ $43.53$ $74.01$ $58.78$ $79.84$
Binder $56.81$ $82.81$ 56.7456.74{56.74}56.74 79.1779.17{79.17}79.17 $50.51$ $78.16$ $55.43$ 81.7281.72{81.72}81.72
ReAcTable 64.1364.13{64.13}64.13 85.7185.71{85.71}85.71 $51.80$ $72.80$ 58.0158.01{58.01}58.01 80.0080.00{80.00}80.00 60.1560.15{60.15}60.15 $81.67$
CoTable $\underline{64.53}$ $\underline{86.22}$ $\underline{59.94}$ $\underline{80.20}$ $\underline{62.22}$ $\underline{85.62}$ $\underline{64.41}$ $\underline{83.20}$
AutoPrep 66.09 87.85 64.75 84.19 63.72 85.72 68.72 85.33
Exp-2: Comparison of AutoPrep with previous SOTA TQA methods with data prep. We compare AutoPrep with TQA methods that consider data prep tasks in their question answering process, using a single LLM agent. The results are reported in Table 3.
The results clearly demonstrate that AutoPrep achieves the new SOTA performance on both the WikiTQ and TabFact datasets across all LLM backbones. In particular, AutoPrep outperforms ICL-Prep with a significant accuracy improvement by 11.00 and 8.33 on WikiTQ and TabFact on average, respectively. Other baselines, except Dater, generally outperform ICL-Prep by introducing specific technical optimizations to data prep. For instance, ReAcTable applies the ReAct paradigm to iteratively generate data prep operations rather than producing all code at once, and it explicitly considers data prep tasks (e.g., filtering and augmentation). Despite this, AutoPrep still outperforms the second best method by considerable improvement. Specifically, when compared with CoTable, AutoPrep achieves better performance, with an average accuracy improvement by 3.05 and 1.96 on WikiTQ and TabFact, respectively.
These improvements are primarily attributed to our multi-agent framework, which effectively tackles the question-aware data prep challenge. The experimental results indicate that data prep is inherently complex and is hard to be resolved by a single, one-size-fits-all solution; instead, a more effective approach is to develop specialized methods for each type. A centralized module can then determine which operations to perform and assign them accordingly. Moreover, AutoPrep covers a broader range of data prep operations than these state-of-the-art TQA methods, thereby addressing gaps (e.g., normalization tasks) not fully explored by previous solutions.
Finding 2: AutoPrep outperforms the the previous SOTA TQA methods with data prep by 3.05 points on WikiTQ and 1.96 points on TabFact, respectively,
In addition, we find that DeepSeek achieves the better overall performance at lower API costs. Thus, we select DeepSeek as our default backbone LLM for subsequent experiments.
@@
Exp-3: Evaluation of AutoPrep on tables with various sizes. To further investigate AutoPrep, we analyze its performance across tables of varying sizes. While all tables in TabFact are small, we categorize the tables in WikiTQ into Small (fewer than 2048 tokens), Medium (2048 to 4096 tokens), and Large (more than 4096 tokens), resulting in 4040 Small tables, 200 Medium tables, and 104 Large tables. The results are reported in Figure 7.
We observe that all baselines exhibit unstable performance, particularly on larger tables. For instance, although CoTable achieves the highest average accuracy among previous SOTA methods, it suffers an accuracy drop of 11.78 when reasoning over large tables. Likewise, ReAcTable shows relatively stable performance on medium tables (dropping by 2.15), yet its performance on large tables remains unsatisfactory (dropping by 17.53).
(a) Small
(b) Medium
(c) Large
Figure 7. Comparison of AutoPrep and other data prep methods on tables with different sizes on the WikiTQ dataset.
In contrast, AutoPrep achieves the highest and most stable performance across tables of varying sizes. When processing medium and large tables, the accuracy of AutoPrep drops by 2.29 and 3.79. The key to its stability is that it employs specialized agents for each data prep task, mitigating the issue of exceeding prompt length limits. Furthermore, each agent generates program-based operators to handle data prep on the entire table, thereby avoiding information loss that occurs when large tables are cropped.
Finding 3: By employing multiple agents and program-based operations, AutoPrep maintains stable performance as table size grows, ensuring that each data prep task is handled effectively without overwhelming a single model.
6.4. Evaluation on Generalization
Exp-4: Evaluating generalization on unseen datasets. As discussed previously, we evaluate the generalization capabilities of AutoPrep on two TabBench datasets. Specifically, we directly use the designed prompting strategies on the WikiTQ dataset, and examine whether these strategies can be generalized to TabBench.
As shown in Figure 8, AutoPrep achieves the highest overall accuracy among all methods. Specifically, compared with the second best method ReAcTable, AutoPrep improves by 5.28, indicating the strong generalization capability of our method. The main reason is that AutoPrep utilizes tool-augmented method for physical plan generation and execution, generalizing well on unseen datasets.
Finding 4: Utilizing tool-augmented execution enhances the generalization capabilities of AutoPrep on unseen datasets.
(a) Numerical Reasoning
(b) Fact Checking
Figure 8. Evaluating Generalization on the TabBench dataset.
6.5. Ablation Studies
@@
Exp-5: Evaluation on the Planner agent. We compare two Planner variants with different high-level operation suggestion methods, namely Direct Prompting and our proposed Chain-of-Clauses method, and report the results in Table 4(a).
@@
We observe that Chain-of-Clauses outperforms Direct Prompting by 7.92, 4.50, 5.27 in accuracy on WikiTQ, TabFact and TabBench respectively. This indicates the superiority of our proposed method in generating more accurate high-level operations. Moreover, we find that the performance improvement on WikiTQ is more significant than that on other datasets. This is because the WikiTQ dataset has relatively large tables and complex questions, which could make the high-level operation suggestion problem more challenging to be solved using Direct Prompting.
Finding 5: Our proposed Chain-of-Clauses method is more effective in suggesting high-level operations.
@@
Figure 9. Two case studies selected from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
Exp-6: Evaluation on the Programmer Agents. We change the low-level operation generation methods in the Programmer agents and keep other settings of AutoPrep as default.
As illustrated in Table 4(b), our proposed tool-augmented method achieves better performance compared with the code generation method. Considering the high-level operations input to the Programmer agents are the same, we can conclude that selecting a function from a function pool and then completing its arguments can generate more accurate and high-quality programs to implement the high-level operations. Moreover, when using demonstrations constructed from same table-question pairs for the prompt of programmer agents, Tool-augmented method can save 18.07% input tokens for low-level operation generation. We also record the error ratio of these two methods, as shown in Table 4(b). The probability of bugs in our tool-augmented method is greatly reduced (e.g., from 4.51% to 1.50%), demonstrating its effectiveness.
Finding 6: Our tool-augmented method performs better in both accuracy and cost, with fewer program errors.
Table 4. Experimental Results of Ablation Studies.
(a) Evaluation on the Planner agent.
Method  WikiTQ  TabFact  TabBench
Direct Prompting $58.17$ $83.35$ $44.83$
Chain-of-Clauses 66.09 87.85 50.10
(b) Evaluation on the Programmer Agents.
Method Metric  WikiTQ  TabFact  TabBench
Code Generation Acc $\uparrow$ $62.82$ $81.97$ $47.26$
Err $\downarrow$ $4.51\%$ $4.50\%$ $4.73\%$
Tool Augmented Acc $\uparrow$ 66.09 87.85 50.10
Err $\downarrow$ 1.50% 0.05% 1.01%
(c) Contribution of Each Programmer Agent in AutoPrep.
Method  WikiTQ  TabFact  TabBench
AutoPrep 66.09 87.85 50.10
- Filter $62.78$ ($-3.31$) $84.98$ ($-2.87$) $47.67$ ($-2.43$)
- Augmenter $61.79$ ${(-4.30)}$ $85.67$ ($-2.18$) $45.44$ ($-4.66$)
- Normalizer $62.02$ ($-4.07$) $83.79$ ${(-4.06)}$ $41.58$ ${(-8.52)}$
Exp-7: Contribution of each agent in AutoPrep. We ablate each Programmer agent including Filter, Augmenter and Normalizer and compare the performance with AutoPrep.
As shown in Table 4(c), for WikiTQ, without column augmentation, the accuracy drops the most (4.30). For TabFact and TabBench, the Normalizer matters the most with an accuracy drop by 4.06 and 8.52. This is because that WikiTQ has more instances requiring string extraction or calculation to generate new columns for answering the question, while normalization is a primary issue in TabFact and TabBench. Moreover, for all datasets, each agent plays an essential role in data preparation for TQA, which brings accuracy improvement by at least 2.43, 2.18 and 4.06.
Finding 7: Solving three types of data preparation tasks in TQA methods in a specialized manner improves the downstream TQA performance.
6.6. Case Studies
@@
This section presents two illustrative examples from TabBench and WikiTQ to qualitatively analyze effectiveness of AutoPrep.
To answer the question in Figure 9(a), CoTable generates an operator chain with an “add_col” operator, which is used to generate a new column Tol based on the division results of two existing columns Expense and Percent. This operator directly prompts an LLM to output a list containing all values of the new column, which is a challenging task for previous LLMs. Thus, CoTable generates a new column with two wrong values which leads to a wrong answer. For ReAcTable, although it generates a logically correct Python program, it ignores the types of existing columns which do not support for numerical calculation. Thus, ReAcTable also fails to augment the Tol column. AutoPrep addresses this issue by first utilizing a Planner agent to generate logical operators specifying data prep requirements, i.e., two $\mathtt{Normalize}$ to normalize column Expense and Percent, a $\mathtt{Augment}$ to generate the overall expenses Tol and a $\mathtt{Filter}$ to select related columns Year and Tol. These logical operators are passed to Programmer agents to generate physical operators consisting of a pre-defined Python function. For example, we call $\mathtt{to\mbox{-}numerical}$ function two times to cast Expense and Percent to numerical values. All physical operators are executed to process the original table and output a prepared table. Finally, an Analyzer agent based on NL2SQL method generates a SQL query to extract the correct answer “4214.17”.
Figure 9(b) illustrates a TQA instance involving large tables and inconsistent values. AutoPrep solves this by splitting the TQA question into two phases, i.e., data prep and data analysis. It first generates logical operators including two $\mathtt{Normalize}$ and one $\mathtt{Filter}$. Next, the Programmer agents implement them with corresponding physical operators, which are executed to produce a prepared table. Based on this, the Analyzer agent extracts the final answer “8”.